A task force of capitail budgeting analysts at Morrison Ltd. collected the following data conceming the drilling and production of known petroleum reserves at an offshore location: Table 6-4 (Use appropriate factor(s) from the table provided. Round the PV factors to 4 decimals.) ST.100,000 Investment in rigging equipment and related personnel costs requred to pump the oll Net increase in inventory and receivables associated with the drilling and production of the reserves. Assume this investment w be recovered at the end of the project 1210,000 Net cash inflow from operations for the expected life of the reserves, by year 2016 2017 2018 2.010,000 3,810,000 1710,000 1.200,000 Salvage value of machinery and equipment at the end of the welrs productive life Cost of capital 4% Required: a. Calculate the net present value of the proposed investment in the driling and production operation. Assume that the investment will be made at the beginning of 2016, and the net cash inflows from operations will be received in a lump sum at the end of each year (Ignore income taxes b. What will the internal rate of retun on this investment be relative to the cost of capital? c. Differences between estmates made by the task force and actual results would have an offect on the actual rate of return on the project. For each estimate, state the offect on the actual ROl if the estimate turns out to be loss than the actual amount finally achieved Estimate Investment cost Actual NPy and ROI will be Annual cash inflows Actual NPV and ROl will be Cost of capital Efect if estimate is less than actual: indic than indicated Actual NPV and ROI wll be
Answer #1
A B C D E F G H I J
2
3 a)
4
5 Net Cash Flow of the project can be represented as below:
6 Year 0 1 2 3
7 Initial investment ($7,100,000)
8 Cash Inflow $2,010,000 $3,610,000 $1,710,000
9 Salvage Value $1,200,000
10 Investment in working capital ($1,210,000)
11 Recovery of working capital $1,210,000
12 Net Cash Flow ($8,310,000) $2,010,000 $3,610,000 $4,120,000 =SUM(G7:G11)
13
14 NPV calculation:
15 NPV of the project is present value of future cash flows discounted at required rate of return less the initial investment.
16 Given the following cash flow and WACC, NPV for the project can be calculated as follows:
17 Year 0 1 2 3
18 Free Cash Flow (FCF) ($8,310,000) $2,010,000 $3,610,000 $4,120,000
19 MARR (i) 4%
20 (P/F,i,n) for each year 0.96 0.92 0.89 =1/((1+$D19)^G17)
21 Present Value of cash flows = FCF*(P/F,i,n) $1,932,692 $3,337,648 $3,662,665 =G18*G20
22 Present value of cash flows $8,933,005.23 =SUM(E21:G21)
23
24 NPV for Project =Present value fo future cash flows - Initial investment
25 $623,005 =D22+D18
26
27 Hence NPV of the Project is $623,005
28
29 b)
30
31 Because NPV of the project is higher than zero, the internal rate of return will be higher than the cost of capital.
32
33 c)
34
35 If investment cost estimate is less than the actual i.e. actually investment cost is higher,
36 therefore the NPV and the ROI will be less than the actual.
37
38 If cash inflow estimate is less than the actual i.e. actually cash inflow is higher,
39 therefore the NPV and the ROI will be higher than the actual.
40
41 If cost of capital estimate is less than the actual i.e. actually cost of capital is higher,
42 therefore the NPV and the ROI will be less than the actual.
43